Q1. from a square sheet of uniform density a portion is removed .find the centre of mass of the remaining portion if the side of the square is A. 

Grade:

4 Answers

shivangi bajpai
11 years ago

simran wd u plz tell me at kind of a portion u mean?

Shiva
6 years ago
Let m1 be the α*a² and m2 be the α*a²/4 where α be the uniform mass density From the question, we have to find the centre of mass of remaining portion. Now we have Xcm = m1*x1 +m2*x2 /m1+m2=( α*a²*0 - α*a²/4*a/2)/α*a²+α*a²= -a³*α/8*4α*a²/5= -a/10There was no possibility to create a diagram
pradyumna Anand
3 years ago
x × mass remaining = mass removed × side/total no. of unshade sides
(m-m/4) × x = (m/4)(a/3)
X=a/3
pradyumna Anand
3 years ago
sorry
(3m/4)x=(m/4)(a/3)
Answer=》x=a/9
In above answer I have done a small error in solving ,this is the correct answer 👌.